OpenStudy (anonymous):

The slope of a line is -7 . The point (10, 3) is on the line. Determine the equation of this line. 10 it should be -7 ovr 10

OpenStudy (anonymous):

If the slope is -7/10: y - 3 = (-7/10)(x - 10) y - 3 = -7/10(x) + 7 y = -7/10(x) + 10